OpenStudy (anonymous):

Cabon-14 has a half life of 5760 years .How old is a fossil that has 1/8 of the substance left?

OpenStudy (dumbcow):

multiply half-life by 3 3*5760 = 17,280

OpenStudy (anonymous):

17,280?

OpenStudy (anonymous):

What grade is this for?

OpenStudy (anonymous):

Algebra 2

OpenStudy (lgbasallote):

shouldnt it be 1440 yrs?

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

are you doing y=a*b^kt or y=a*e^kt?

OpenStudy (lgbasallote):

well 1/8 is 1/4 of 1/2 so i divided 5760 by 4...dunno if it should be done that way...is it?

OpenStudy (dumbcow):

usually for carbon half-life you use y = a*e^kt 1/2 = e^5760k 1/2 ^3 = 1/8 --> 1/8 = e^3(5760k) thus t = 3*5760

OpenStudy (anonymous):

\[y=A_0 e^{kt}\] where A_0 is the initial amount. when y=0.5A_0, t=5760 \[0.5A_0=A_0 e^{5760k}\]\[0.5=e^{5760k}\]\[ln0.5=lne^{5760k}\]\[-0.693=5760k\]\[k= -0.00012 \] when y=1/8 A_0 \[A_0/8=A_0 e^{-0.00012t}\]\[1/8=e^{-0.00012t}\]\[ln1/8=lne^{5760k}\]\[-2.079=-0.00012t\]\[t= 17 328 \]

OpenStudy (anonymous):

TY